ACJ is a company that specializes in business continuity auditing for businesses of all sizes. At this time, the company is planning to bid on two opportunities (projects) sponsored by the Small Business Administration (SBA), to audit how well prepared are small businesses to confront risky events in several counties throughout the United States. All interested companies could bid to audit businesses in more than one county. ACJ is interested in a county in Florida and another in Arizona. The former has 25 small businesses and the latter 35. The winner of each bid will receive $375,000 for auditing all businesses of the county, regardless of its size or complexity. (Note: For example, a company that wins three bids would get a total of $1,125,000, and a company that wins four bids would get $1,500,000.)
Based on the company’s experience auditing in Florida, management estimates that the first audit will take 200 labor hours and the second 180 labor hours. There is a fixed cost of $25,000 for operating in Florida and a standard rate of $35/hour. In Arizona, the second audit will take 160 labor hours and the fourth 136 labor hours. The fixed cost in Arizona is $20,000 and the standard wage rate is $25/hour. A company’s policy requires that only a program that returns at least 2.5 times its total cost should be engaged. (Note: In this case the return of a program will depend on the number of awarded bids.)
Before submitting the bid, ACJ’s management would like your group to decide whether they should bid for auditing both counties or only one. (Explain and illustrate.)
Answer #1

Cost estimate foe Florida:

Total number of labor hours = 200+180 = 380

Standard labor rate per hour = $35

Total labor cost per business= 35*380 = $13300

Number of small businesses in Florida = 25

Total labor cost = 25*13300 = $332500

Fixed cost = $25000

Total cost of operating in Florida = 332500+25000 = $357500